# Identifiability of nonlinear ODE Models with Time-Varying Parameters: the General Analytical Solution and Applications in Viral Dynamics

### v4

@@ 3.3.1 $\boldsymbol{{}^{m}\mu}$ and $\boldsymbol{{}^{m}\nu}$
Note that the entries with $i,~{}j=1,\ldots,m$ are the same entries of the matrix that consists of the first $m$ columns of the unknown input reconstructability matrix from $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$. By construction, this submatrix is full rank. As a result, the tensor ${}^{m}\mu$ is also non singular and we denote by ${}^{m}\nu$ its inverse.
3.3.2 $\boldsymbol{{}^{m}\widehat{g}^{\alpha}}$
@@
Starting from ${}^{m}\nu$ we build the new vector fields ${}^{m}\widehat{g}^{0},\ldots,^{m}\widehat{g}^{m}$ defined as follows:
$^{m}\widehat{g}^{\alpha}=\sum_{\beta=0}^{m}~{}^{m}\nu^{\alpha}_{\beta}~{}g^{\beta},~{}~{}\alpha=0,1,\ldots,m.$ (3.7)
Once all the quantities ${}^{m}\mu$, ${}^{m}\nu$, and ${}^{m}\widehat{g}^{\alpha}$ are computed, the loop at lines 9-14 of Algorithm 1 ends by computing
$\Omega=\left<\left.f^{1},\ldots,f^{m_{u}},\left\{\begin{array}[]{lll}{}^{m}\widehat{g}^{0},&\ldots,&{}^{m}\widehat{g}^{m}\\ g^{m+1},&\ldots,&g^{m_{w}}\\ \end{array}\right\}~{}\right|\Omega\right>.$
(note that, for TV systems, the Lie derivative operator along ${}^{m}\widehat{g}^{0}$, i.e., $\mathcal{L}_{{}^{m}\widehat{g}^{0}}$, must be replaced by .$\accentset{\mbox{\large.}}{\mathcal{L}}_{{}^{m}\widehat{g}^{0}}:=\mathcal{L}_{{}^{m}\widehat{g}^{0}}+\frac{\partial}{\partial t}$). The loop is executed when the conditions given at Line 9 are satisfied. Note that, as $\textnormal{deg}_{w}\left(\Omega\right)$ cannot exceed $m_{w}$, the loop cannot be repeated more than $m_{w}-1$ times.
After the execution of this loop, Algorithm 1 computes the integer $\widehat{k}^{m}=s^{m}_{x}+r^{m}$, where the integers $r^{m}$ and $s^{m}_{x}$ are defined similarly to the integers $r$ and $s$ introduced in Section 3.2.3:
  • $s^{m}_{x}$ is the smallest integer such that ${}^{x}\Omega^{m}_{s^{m}_{x}}=~{}^{x}\Omega^{m}_{s^{m}_{x}-1}$, where ${}^{x}\Omega^{m}_{k}$ is the codistribution at the $k^{th}$ step computed by Algorithm 2. Note that $s^{m}_{x}\leq n-m+1$.
  • $r^{m}$ is the smallest integer such that ${}^{m}\Delta_{r^{m}+1}=~{}^{m}\Delta_{r^{m}}$, where ${}^{m}\Delta_{k}$ is the distribution at the $k^{th}$ step computed by Algorithm 3. Note that $r^{m}\leq n-1$ (it is even $r^{m}\leq n-\textnormal{dim}\left\{\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}\right\}$).
We have $\widehat{k}^{m}=s^{m}_{x}+r^{m}\leq 2n-m+1-\textnormal{dim}\left\{\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}\right\}\leq 2n-m$.
Below, we provide a description of Algorithms 2 and 3.
@@
It computes the distribution ${}^{m}\Delta$. The initialization step sets the distribution equal to the span of the vector fields $f^{1},\ldots,f^{m_{u}}$. Then, the recursive step adds to ${}^{m}\Delta_{k-1}$ the term $\sum_{\beta=0}^{m_{w}}\left[~{}^{m}\Delta_{k-1}\right]^{\beta}$. The convergence is attained at the smallest integer $j$ such that ${}^{m}\Delta_{j}=~{}^{m}\Delta_{j-1}$ and $j\leq n-\textnormal{dim}\left\{\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}\right\}+1$. We set $r^{m}+1=j$. Note that, because of the presence of the state augmentation in the loop (Line 5 of Algorithm 3), proving the validity of the above convergence property is non trivial (see Section 6.2.1 of [2] and in particular the proof of Proposition 1 given in Appendix D of [2]).
Note that, the system augmentation performed by the $\mathcal{A}$ operation executed by Algorithms 2 and 3, does not reset the system in Algorithm 1. The system defined in Algorithm 1 can only be reset (augmented) by the operation $\mathcal{A}$ that is executed at Line 26 of Algorithm 1. The execution of Algorithm 2 and Algorithm 3 only returns the integer $\widehat{k}^{m}=s^{m}_{x}+r^{m}$. All remaining quantities remain unchanged.
3.3.3 The nested loop
Algorithm 1 continues by computing the codistributions $\Omega_{k}$, $k=0,1,\ldots$. For positive $k$, this is obtained by executing the nested loop (Lines 24-42).  Each iteration, first augments the system (Line 26 of Algorithm 1). Note that, this operation updates the set of vector fields $g^{0},~{}g^{1},\ldots,~{}g^{m}$, in accordance with (3.2). As a result, it is necessary to recompute ${}^{m}\widehat{g}^{0},~{}^{m}\widehat{g}^{1},\ldots,~{}^{m}\widehat{g}^{m}$ by using (3.7). This is carried out at the next line (Line 27). Then, Algorithm 1 computes $\Omega_{k}$ (Lines 29-32). The last part of the loop contains two if statements (Line 33, and Line 38). Each of them, when the condition is met, interrupts the loop. In [2], we proved that at least one of these two conditions is met in at most
$\widehat{k}^{m}+n-\dim\left(\Omega_{0}\right)\leq\widehat{k}^{m}+n-m\leq 3n-2m$
steps, where $n$ is the dimension of the original state, i.e., the state of the system before entering the nested loop. Note that, in [2] (Section 7), we actually proved that one of the two if conditions is met in at most $n-m$. This because the first $\widehat{k}^{m}$ steps were included in the initialization step (this is less efficient because the condition of the second if can be met during the execution of these first $\widehat{k}^{m}$ steps).
When the condition of the first if statement (Line 33) is met, not only the loop is interrupted but also the execution of Algorithm 1 terminates. In this case, the system is not canonic with respect to its unknown inputs. Algorithm 1 provides the entire observable codistribution. In addition, the augmented system is the one with the highest unknown input degree of reconstructability (see Definition 3.1). When the loop is interrupted by the second if statement (Line 38), the execution of the nested loop is interrupted.  Again, as $m\leq m_{w}$, the condition at Line 38 cannot be met more than $m_{w}-1$ times.
@@ 3.4 Illustration by the case study
and $h_{1}=\phi-\theta$. We run Algorithm 1. Line 2 provides $\Omega=\textnormal{span}\{\nabla h_{1}\}=\textnormal{span}\{[0,1,-1]\}$, and $\textnormal{deg}_{w}(\Omega)=1$, as the unknown input observability matrix from $h_{1}$ is the $1\times 1$ matrix equal to $\mathcal{L}_{g^{1}}h_{1}=-1\neq 0$ (its rank is 1). The system is in canonical form with respect to its unknown input. Line 45 trivially selects $\widetilde{h}_{1}=h_{1}$. Then, Line 46 provides:
@@
The final step of Algorithm 1 is the execution of Line 47, namely:
@@
We run Algorithm 1. Line 2 provides $\Omega=\textnormal{span}\{\nabla h_{1}\}=\textnormal{span}\{[0,1,-1]\}$, and $\textnormal{deg}_{w}(\Omega)=1$, as the unknown input observability matrix from $h_{1}$ is the $1\times 1$ matrix equal to $\mathcal{L}_{g^{1}}h_{1}=\frac{\sin(\theta-\phi)}{\rho}\neq 0$ (its rank is 1). The system is in canonical form with respect to its unknown input. Line 45 trivially selects $\widetilde{h}_{1}=h_{1}$. Then, Line 46 provides:
@@
Finally, Line 46 computes $\widetilde{\mathcal{O}}$. From (3.5), by a direct computation we obtain:
@@
The final step of Algorithm 1 is the execution of Line 47, namely:
@@ 4.3 Identifiability for any system
We start by running Algorithm 1. By doing this, we automatically obtain the observability codistribution, $\mathcal{O}$. In addition, the algorithm provides a new system, which is an unknown input extension of the original system with the highest unknown input degree of recostructability. From now on, we refer to this system. Its structure still satisfies (2.2). We denote the dimension of its state by $n$ and we denote its unknown input degree of recostructability (which is the highest) by $m$. Note that, a set of $m$ observable functions, $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$, is also available, after the execution of Algorithm 1. The new system could coincide with the original system, meaning that the UI degree of reconstructability of the original system is already the highest. In this case, Algorithm 1 returns the original system where only the order of the unknown inputs could have changed. Indeed, when $m<m_{w}$, the system returned by Algorithm 1 is ordered in accordance with the $\mathcal{R}$ operation, defined in Section 3.1.3. In other words, the rank of the first $m$ columns of $\mathcal{RM}\left(\widetilde{h}_{1},\ldots,\widetilde{h}_{m}\right)$ is $m$.
@@ 4.3.1 Unknown input symmetries due to the system non canonicity
Let us consider the system returned by Algorithm 1 and let us suppose that $m<m_{w}$ (system not canonic with respect to the unknown inputs). For this system, the following $m_{w}-m$ column vectors:
@@ 4.3.2 Possible unknown input symmetries due to the state unobservability
Let us consider the system returned by Algorithm 1, and let us denote by $\xi\in\mathcal{O}^{\bot}$ a given symmetry of the state. The $m_{w}-$ column vector ${}^{u}\chi$, with components:
$\left\{\begin{array}[]{ll}{}^{u}\chi_{k}=-\sum_{i=1}^{m}~{}^{m}\nu^{i}_{k}\left(\xi^{0}_{i}+\sum_{j=1}^{m}\xi^{j}_{i}w_{j}\right)&k=1,\ldots,m\\ {}^{u}\chi_{k}=0~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}k=m+1,\ldots,m_{w}\\ \end{array}\right.$ (4.6)
@@
$\xi^{\alpha}_{i}=\left\{\begin{array}[]{ll}\left(\frac{\partial}{\partial x}\left(\mathcal{L}_{g^{0}}\widetilde{h}_{i}+\frac{\partial\widetilde{h}_{i}}{\partial t}\right)\right)\cdot\xi&\alpha=0\\ \left(\frac{\partial}{\partial x}\mathcal{L}_{g^{\alpha}}\widetilde{h}_{i}\right)\cdot\xi&\alpha=1,\ldots,m\\ \end{array}\right.$ (4.7)
@@
When the system is canonic with respect to its unknown inputs ($m=m_{w}$), the unknown input symmetry becomes:
${}^{u}\chi_{k}=-\sum_{i=1}^{m_{w}}~{}\nu^{i}_{k}\left(\xi^{0}_{i}+\sum_{j=1}^{m_{w}}\xi^{j}_{i}w_{j}\right),~{}~{}~{}~{}~{}k=1,\ldots,m_{w}.$
@@ 4.3.3 The systematic procedure
The systematic procedure to automatically obtain the identifiability of all the time-varying parameters (or the reconstructability of all the unknown inputs) is Algorithm 4. It is based on the results stated by Theorems 4.2 and 4.3. It starts by running Algorithm 1 on the system under investigation (Line 1). This provides the observability codistribution $\mathcal{O}$, the highest UI degree of reconstructability with respect to the unknown inputs, $m$, and a new system $\Sigma^{\prime}$ which is an unknown input extension of the original system with the highest UI degree of reconstructability. Note that, the identifiability analysis regards the new system $\Sigma^{\prime}$. Hence, we investigate the reconstructability of the unknown inputs of this system. These unknown inputs may differ from the original ones for the following two reasons:
@@
Note that, in any case, we also obtain the reconstructability properties of the original unknown inputs. Indeed, if a given UI of $\Sigma^{\prime}$ is the time derivative of a given UI of the original system, it certainly means that the original UI was included in the state. To check its identifiability it suffices to verify if its differential belongs to $\mathcal{O}$. Algorithm 4 will tell us if the corresponding UI of $\Sigma^{\prime}$ is identifiable (i.e., if the time derivative of a given order of the original UI is identifiable).
@@
Lines 5 and 6 compute the orthogonal distribution $\mathcal{O}^{\bot}$ and a basis of it. In particular, its generators are $\xi^{1},\ldots,\xi^{S}$. Then, the loop at Lines 7-9 computes the corresponding symmetries of the unknown input vector, i.e., ${}^{u}\chi^{1},\ldots,~{}^{u}\chi^{S}$, on the basis of Theorem 4.3.
@@
Finally, the loop at Lines 10-16 checks the reconstructability of all the unknown inputs $w_{1},\ldots,w_{m_{w}}$. Specifically, the $j^{th}$ unknown input ($w_{j}$) can be reconstructed if and only if:
@@
Note that, when $S=0$ (and this is the case when the state that characterizes $\Sigma^{\prime}$ is observable) the first $m$ unknown inputs, $w_{1},\ldots,w_{m}$, can be reconstructed. If in addition the system is canonic with respect to its unknown inputs, all the unknown inputs can be reconstructed. Note that the observability of the state is a sufficient (but not a necessary) condition to have the reconstructability of the first $m$ unknown inputs. Indeed, there are cases where the state is unobservable (i.e., $S>0$) but all the symmetries ${}^{u}\chi^{1},\ldots,~{}^{u}\chi^{S}$ vanish, simultaneously (e.g., this occurs in the first scenario of our case study, as discussed in Section 4.4.3).
1:  Run Algorithm 1 and compute $\mathcal{O}$, $m$, $\Sigma^{\prime}$
@@
4:  end if
@@
7:  for $i_{s}=1:S$ do
8:     Set $\xi=\xi^{i_{s}}$, and compute the corresponding ${}^{u}\chi$ (from (4.6) and (4.7)), and denote it by ${}^{u}\chi^{i_{s}}$.
9:  end for
10:  for $j=1:m_{w}$ do
11:     if ${}^{c}\chi^{1}_{j}=\ldots=~{}^{c}\chi^{m_{w}-m}_{j}=~{}^{u}\chi^{1}_{j}=\ldots=~{}^{u}\chi^{S}_{j}=0$ then
12:        $w_{j}$ can be reconstructed
13:     else
14:        $w_{j}$ cannot be reconstructed
15:     end if
16:  end for

### v5

@@ 3.3.1 $\boldsymbol{\mu}$ and $\boldsymbol{\nu}$
Note that the entries with $i,~{}j=1,\ldots,m$ are the same entries of the matrix that consists of the first $m$ columns of the unknown input reconstructability matrix from $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$. By construction, this submatrix is full rank. As a result, the tensor $\mu$ is also non singular and we denote by $\nu$ its inverse.
3.3.2 $\boldsymbol{\widehat{g}^{\alpha}}$
@@
Starting from $\nu$ we build the new vector fields $\widehat{g}^{0},\ldots,\widehat{g}^{m}$ and $\widehat{g}^{m+1},\ldots,\widehat{g}^{m_{w}}$. They are defined as follows:
$\widehat{g}^{\alpha}=\sum_{\beta=0}^{m}~{}\nu^{\alpha}_{\beta}~{}g^{\beta},~{}~{}\alpha=0,1,\ldots,m,~{}~{}~{}~{}\widehat{g}^{k}:=g^{k}-\sum_{\alpha=0}^{m}\widehat{g}^{\alpha}\mathcal{L}_{g^{k}}\widetilde{h}_{\alpha},~{}~{}k=m+1,\ldots,m_{w}$ (3.7)
3.3.3 Algorithm 2
This algorithm returns the Boolean Finish. When Finish==True, it means that $\Omega_{*}$ is the observability codistribution. When Finish==False, the execution of the algorithm only tells us that there exists an observable function that increases the unknown inputs degree of reconstructability. Let us denote this function by $\theta$. Algorithm 2 does not provide $\theta$. $\theta$ is determined by the remaining steps of the iteration of the main loop of Algorithm 1, under execution. Note that, when Finish==False, Algorithm 2 returns $\Sigma_{*}$, which can differ from $\Sigma$ and is the suitable system to compute $\theta$. In addition, when Finish==False, Algorithm 2 provides $k_{*}$, $i_{*}$, and $q_{*}$, which are used in the main loop of Algorithm 1 to compute $\theta$.
3.3.4 The $\boldsymbol{\widehat{g}}$ vector field
It appears at Line 13 of Algorithm 1. It is defined as follows:
${}\widehat{g}\triangleq\sum_{\beta=0}^{m}~{}\widehat{g}^{\beta}~{}v_{\beta}$ (3.8)
with ($\beta=0,1,\ldots,m$)
$v_{\beta}\triangleq\left[\begin{array}[]{ll}\sum_{\gamma=0}^{m}~{}\mu^{\gamma}_{\beta}~{}w_{\gamma}+\sum_{k=m+1}^{m_{w}}(\mathcal{L}_{g^{k}}\widetilde{h}_{\beta})~{}w_{k}&k_{*}=1\\ \sum_{\gamma=0}^{m}~{}\mu^{\gamma}_{\beta}~{}w_{\gamma}&k_{*}>1\\ \end{array}\right.$ (3.9)
[…]
The expression of $\psi_{k}^{i}$ is obtained automatically and recursively. We have:
$\psi_{0}^{i}=f^{i},~{}~{}~{}~{}~{}\psi_{k+1}^{i}=\sum_{\gamma=0}^{m}[g^{\gamma},~{}\psi^{i}_{q}]w_{\gamma}+\sum_{\alpha=0}^{m}\sum_{i_{\alpha}=0}^{k}\frac{\partial\psi_{k}^{i}}{\partial w_{\alpha}^{(i_{\alpha})}}w_{\alpha}^{(i_{\alpha}+1)}.$ (3.11)
For time-varying systems $\psi_{k+1}^{i}\rightarrow\psi_{k+1}^{i}+\frac{\partial\psi_{k}^{i}}{\partial t}$.
Once determined, we must eliminate $w_{\alpha}$, $w_{\alpha}^{(1)}$, $\ldots$, $w_{\alpha}^{(i_{\alpha}+1)}$ by expressing them in terms of $v_{\alpha}$, $v_{\alpha}^{(1)}$, $\ldots$, $v_{\alpha}^{(i_{\alpha}+1)}$. This is achieved by using the inverse of Equation (3.9) above, i.e., $w_{\alpha}=\sum_{\beta=0}^{m}~{}\nu_{\alpha}^{\beta}v_{\beta}$.
@@ 3.3.4 The $\boldsymbol{\widehat{g}}$ vector field
It computes the distribution $\Delta$. The initialization step sets the distribution equal to the span of the vector fields $f^{1},\ldots,f^{m_{u}}$. Then, the recursive step adds to $\Delta_{k-1}$ the term $\sum_{\beta=0}^{m_{w}}\left[\Delta_{k-1}\right]^{\beta}$. The convergence is attained at the smallest integer $j$ such that $\Delta_{j}=\Delta_{j-1}$ and $j\leq n-\textnormal{dim}\left\{\textnormal{span}\left\{f^{1},\ldots,f^{m_{u}}\right\}\right\}+1$. We set $r^{m}+1=j$. Note that, because of the presence of the state augmentation in the loop (Line 5 of Algorithm 4), proving the validity of the above convergence property is non trivial (see Section 6.2.1 of [2] and in particular the proof of Proposition 1 given in Appendix D of [2]).
Note that, the system augmentation performed by the $\mathcal{A}$ operation executed by Algorithms 3 and 4, does not reset the system in Algorithm 1. The execution of Algorithm 3 and Algorithm 4 only returns the integer $\widehat{k}^{m}=s^{m}_{x}+r^{m}$. All remaining quantities remain unchanged.
@@ 3.4 Illustration by the case study
and $h_{1}=\phi-\theta$. We run Algorithm 1. Line 1 provides $\Omega=\textnormal{span}\{\nabla h_{1}\}=\textnormal{span}\{[0,1,-1]\}$, and $\textnormal{deg}_{w}(\Omega)=1$, as the unknown input observability matrix from $h_{1}$ is the $1\times 1$ matrix equal to $\mathcal{L}_{g^{1}}h_{1}=-1\neq 0$ (its rank is 1). The system is in canonical form with respect to its unknown input. Line 18 trivially selects $\widetilde{h}_{1}=h_{1}$. Then, Line 19 provides:
@@
The final step of Algorithm 1 is the execution of Line 20, namely:
@@
We run Algorithm 1. Line 1 provides $\Omega=\textnormal{span}\{\nabla h_{1}\}=\textnormal{span}\{[0,1,-1]\}$, and $\textnormal{deg}_{w}(\Omega)=1$, as the unknown input observability matrix from $h_{1}$ is the $1\times 1$ matrix equal to $\mathcal{L}_{g^{1}}h_{1}=\frac{\sin(\theta-\phi)}{\rho}\neq 0$ (its rank is 1). The system is in canonical form with respect to its unknown input. Line 17 trivially selects $\widetilde{h}_{1}=h_{1}$. Then, Line 19 provides:
@@
From (3.5), by a direct computation we obtain:
@@
The final step of Algorithm 1 is the execution of Line 20, namely:
@@ 4.3 Identifiability for any system
We start by running Algorithm 1. By doing this, we automatically obtain the observability codistribution, $\mathcal{O}$. In addition, the algorithm provides a new system, which is an unknown input extension of the original system with the highest unknown input degree of recostructability. Note that, the unknown input extension returned by Algorithm 1 does not guarantee that the following condition is met:
$\mathcal{L}_{g^{k}}\widetilde{h}_{i}=0,~{}~{}\forall i=1,\ldots,m,~{}~{}\forall k=m+1,\ldots,m_{w}.$ (4.4)
On the other hand, when for a given $k\in[m+1,\ldots,m_{w}]$ there exists $i\in[1,\ldots,m]$ such that $\mathcal{L}_{g^{k}}\widetilde{h}_{i}\neq 0$, it suffices to consider the unknown input extension that also includes $w_{k}$.
Hence, from now on we consider the unknown input extension where the condition in (4.4) is honoured. From now on, we refer to this system. Its structure still satisfies (2.2). We denote the dimension of its state by $n$ and we denote its unknown input degree of recostructability (which is the highest) by $m$. Note that, a set of $m$ observable functions, $\widetilde{h}_{1},\ldots,\widetilde{h}_{m}$, is also available, after the execution of Algorithm 1. The new system could coincide with the original system, meaning that the UI degree of reconstructability of the original system is already the highest. In this case, Algorithm 1 returns the original system where only the order of the unknown inputs could have changed. Indeed, when $m<m_{w}$, the system returned by Algorithm 1 is ordered in accordance with the $\mathcal{R}$ operation, defined in Section 3.1.3. In other words, the rank of the first $m$ columns of $\mathcal{RM}\left(\widetilde{h}_{1},\ldots,\widetilde{h}_{m}\right)$ is $m$.
@@ 4.3.1 Unknown input symmetries due to the system non canonicity
Let us consider the unknown input extension where the unknown input degree of reconstructability coincides with its highest unknown input degree of reconstructability and that honours the condition in (4.4) [^3] This is the system returned by Algorithm 1 and, if the condition in (4.4) is not honoured by a given $k$, it suffices to also include $w_{k}$ in the state, as explained above.. Let us suppose that $m<m_{w}$ (system not canonic with respect to the unknown inputs). For this system, the following $m_{w}-m$ column vectors:
@@ 4.3.2 Possible unknown input symmetries due to the state unobservability
Let us consider the unknown input extension where the unknown input degree of reconstructability coincides with its highest unknown input degree of reconstructability and that honours the condition in (4.4) [^4] This is the system returned by Algorithm 1 and, if the condition in (4.4) is not honoured by a given $k$, it suffices to also include $w_{k}$ in the state, as explained above.. Let us denote by $\xi\in\mathcal{O}^{\bot}$ a given symmetry of the state. The $m_{w}-$ column vector ${}^{u}\chi$, with components:
$\left\{\begin{array}[]{ll}{}^{u}\chi_{k}=-\sum_{i=1}^{m}\nu^{i}_{k}\left(\xi^{0}_{i}+\sum_{j=1}^{m}\xi^{j}_{i}w_{j}\right)&k=1,\ldots,m\\ {}^{u}\chi_{k}=0~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}k=m+1,\ldots,m_{w}\\ \end{array}\right.$ (4.7)
@@
$\xi^{\alpha}_{i}=\left\{\begin{array}[]{ll}\mathcal{L}_{\xi}\left(\mathcal{L}_{g^{0}}\widetilde{h}_{i}+\frac{\partial\widetilde{h}_{i}}{\partial t}\right)&\alpha=0\\ \mathcal{L}_{\xi}\left(\mathcal{L}_{g^{\alpha}}\widetilde{h}_{i}\right)&\alpha=1,\ldots,m\\ \end{array}\right.$ (4.8)
@@ 4.3.3 The systematic procedure
The systematic procedure to automatically obtain the identifiability of all the time-varying parameters (or the reconstructability of all the unknown inputs) is Algorithm 5. It is based on the results stated by Theorems 4.2 and 4.3. It starts by running Algorithm 1 on the system under investigation (Line 1). This provides the observability codistribution $\mathcal{O}$, the highest UI degree of reconstructability with respect to the unknown inputs, $m$, and a new system $\mathcal{E}$ which is an unknown input extension of the original system with the highest UI degree of reconstructability. Note that, the identifiability analysis regards the new system $\mathcal{E}$. Hence, we investigate the reconstructability of the unknown inputs of this system. These unknown inputs may differ from the original ones for the following two reasons:
@@
Note that, in any case, we also obtain the reconstructability properties of the original unknown inputs. Indeed, if a given UI of $\mathcal{E}$ is the time derivative of a given UI of the original system, it certainly means that the original UI was included in the state. To check its identifiability it suffices to verify if its differential belongs to $\mathcal{O}$. Algorithm 5 will tell us if the corresponding UI of $\mathcal{E}$ is identifiable (i.e., if the time derivative of a given order of the original UI is identifiable).
@@
Lines 5 and 6 compute the orthogonal distribution $\mathcal{O}^{\bot}$ and a basis of it. In particular, its generators are $\xi^{1},\ldots,\xi^{S}$. Then, the loop at Lines 8-10 computes the corresponding symmetries of the unknown input vector, i.e., ${}^{u}\chi^{1},\ldots,~{}^{u}\chi^{S}$, on the basis of Theorem 4.3.
@@
Finally, the loop at Lines 12-16 checks the reconstructability of all the unknown inputs $w_{1},\ldots,w_{m_{w}}$. Specifically, the $j^{th}$ unknown input ($w_{j}$) can be reconstructed if and only if:
@@
Note that, when $S=0$ (and this is the case when the state that characterizes $\mathcal{E}$ is observable) the first $m$ unknown inputs, $w_{1},\ldots,w_{m}$, can be reconstructed. If in addition the system is canonic with respect to its unknown inputs, all the unknown inputs can be reconstructed. Note that the observability of the state is a sufficient (but not a necessary) condition to have the reconstructability of the first $m$ unknown inputs. Indeed, there are cases where the state is unobservable (i.e., $S>0$) but all the symmetries ${}^{u}\chi^{1},\ldots,~{}^{u}\chi^{S}$ vanish, simultaneously (e.g., this occurs in the first scenario of our case study, as discussed in Section 4.4.3).
1:Run Algorithm 1 and compute $\mathcal{O}$, $m$, $\mathcal{E}$. In the case, extends $\mathcal{E}$ by including $w_{k}$ to honour (4.4).
@@
4:     Set ${}^{c}\chi^{i}=\left[\begin{array}[]{l}0_{m}\\ e_{i}\\ \end{array}\right],~{}~{}~{}i=1,\ldots,m_{w}-m$, as in (4.6).
@@
7:Set $\Delta^{u}=\left\{\right\}$
8:for $i_{s}=1:S$ do
9:     Set $\xi=\xi^{i_{s}}$, and compute the corresponding ${}^{u}\chi$ (from (4.7) and (4.8)), and denote it by ${}^{u}\chi^{i_{s}}$.
[…]
12:for $j=1:m_{w}$ do
13:     if ${}^{c}\chi^{1}_{j}=\ldots=~{}^{c}\chi^{m_{w}-m}_{j}=~{}^{u}\chi^{1}_{j}=\ldots=~{}^{u}\chi^{S}_{j}=0$ then
14:         $w_{j}$ can be reconstructed
15:     else
16:         $w_{j}$ cannot be reconstructed